# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ 6.2 Evaluation of the Encoding-Decoding Direction Pairs
In this experiment, signal vectors are jointly learned with the concept detectors in an unsupervised fashion, employing every proposed facet of the method as laid out in Section 4. Concept detectors are evaluated for their ability to detect each one of the ground-truth concepts, while signal vectors are evaluated for their alignment with the ground-truth concept encoding directions.
Decoding Directions (Concept Detectors): Using the synthetic pixel dataset $\{{\bm{x}}_{{\bm{p}}_{1}}\}\cup\{{\bm{x}}_{{\bm{p}}_{2}}\}$, we assess the ability of each learned concept detector in detecting each one of the ground-truth concepts. Since the method is unsupervised, the name of the concept that each detector detects is unknown. For this reason, we need to implement direction labeling (Section 3.4). In Table 1 we present the Intersection over Union scores for each detector against actual concept classes. Zero values indicate complete purity and no mixing of the concepts, while scores of 1.0 indicate perfect concept detection. As an additional step, we also examine how well the learned filters extract signal values from representations. Since distractors and the latent space bias $\bm{c}$} are not directly estimated by our method, we can only quantify the extracted signal values as a deviation from the dataset’s average value (See also (16) and Section A.5). The Root Mean Squared Error (RMSE) between these extracted values and the ground truth, after subtracting the mean signal value, is noted as $0.06$. We also verified that in this experiment where the distractor components are independent of the concept content contained in the features, the learned filters were orthogonal to all the vectors in the distractor basis $\bm{D}$}, without explicitly enforcing this to happen.
Encoding Directions: Fig. 8 - orange indicates perfect alignment between the learned signal vectors and the ground-truth concept encoding directions. We stress the fact that in this experiment signal vectors were learned in an unsupervised manner together with the concept detectors, and this is different from the approach we took in Section 6.1, where we evaluated signal vectors as estimators under ideal input conditions.
@@ 6.3 Ablation Study
In this synthetic experiment we also consider a case study where we learn the direction pairs by omitting $\mathcal{L}^{fso}$ from the objective function, i.e. learning the direction pairs without (/wo) it. In that case, the concept detectors still discovered the ground-truth directional clusters, as in Table 1. However, the decoding directions were less reliable in extracting the ground-truth signal values. The RMSE between the extracted values and the ground truth, after subtracting the mean signal value, was found to be $0.31$, which is substantially inferior to the standard case discussed above. The latter has an impact on the fidelity of the signal vectors. Fig. 8 purple visualizes the cosine similarity between the learned signal vectors and the ground-truth encoding directions for this case. The experiment shows that the signal direction estimation is less effective compared to when $\mathcal{L}^{fso}$ is taken into account.
6.4 Evaluation against other State-of-the-Art
The ground-truth encoding directions of this example (Table 14 in Appendix A.4) contain both positive and negative components, and the relationship among the encoding directions (and the distractors) is in general not orthogonal. In theory, and without the need for practical experiments, this example cannot be addressed by NMF, K-Means, or PCA. NMF would produce a signal basis of non-negative components. Similarly, the cluster centers of K-Means would point toward the positive quartile where the centroids reside. Finally, the non-orthogonal nature of the ground truth encoding directions implies that the PCA’s solution space is insufficient.
@@ 7 Experiments on Deep Image Classifiers
In this section we apply our method in real-world scenarios. We perform a multifaceted analysis with the aim to shed light on every merit of the method, the method’s relation to other supervised and unsupervised direction learning approaches, and finally its utility in real-world applications. In the following sections we a) assess the faithfulness of the learned encoding-decoding direction pairs b) provide qualitative segmentation results obtained by utilizing our learned concept detectors c) compare with the unsupervised state-of-the-art in interpretability and influence terms d) conduct an ablation study to quantify the contribution of each proposed component e) compare with supervised direction learning with the aim to provide more insights on the interpretability and influence of the directions f) provide application use cases in which the learned direction pairs assist in providing global, local and counterfactual model explanations g) provide an example application in model correction.
@@ 7.1 Direction Learning and Evaluation Protocol
We apply our Encoding-Decoding Direction Pairs (EDDP) on the last convolution layer of five different networks: ResNet18 (He et al., 2016) trained on Places365 (Zhou et al., 2017), ResNet50 (He et al., 2016) trained on Moments in Time (MiT) (Monfort et al., 2019) as well as EfficientNet (b0) (Tan & Le, 2019), Inception-v3 (Szegedy et al., 2016) and VGG16 (Simonyan & Zisserman, 2014) trained on ImageNet (Deng et al., 2009). For learning the direction pairs and quantitatively assessing their interpretability and their influence on network predictions, we expand on the protocol introduced in Doumanoglou et al. (2023). For all the networks that we study, when learning the direction pairs we use the Broden (Bau et al., 2017) concept dataset, except for ResNet50 for which we use Broden-Action Ramakrishnan et al. (2019). The concept datasets feature dense pixel annotations; Broden includes $1197$ concepts across 63K~63K images in $5$ concept categories (object, part, material, texture, color), while Broden-Action incorporates an additional action category with $210$ labels and 23K~23K more images. We emphasize that when learning the direction pairs we do not make use of the annotations that complement the concept datasets.
@@
For baselines, we consider directional clusterings based on either PCA, NMF, or the natural latent space basis. For our EDDP, we examine two variants, namely EDDP-U and EDDP-C. These are differentiated based on the choice of feature manipulation strategy used in Uncertainty Region Alignment, with EDDP-U referring to UFM and EDDP-C to CFM. Unless stated otherwise, our Encoding-Decoding Direction Pairs use the Augmented Lagrangian Loss from Section 4.5, all losses from Sections 4.2 and 3.3 (except $\mathcal{L}^{s}$), along with either $\mathcal{L}^{uur}$ (EDDP-U) or $\mathcal{L}^{cur}$ and $\mathcal{L}^{fso}$ (EDDP-C) from Sections 4.3 and 4.4. Other method parameters for each case are detailed in Section A.7. We emphasize that EDDP-U does not include the use of $\mathcal{L}^{fso}$ to learn the directions. We remind that the main purpose of $\mathcal{L}^{fso}$ is to contribute towards a faithful estimation of the encoding directions. While this was critical for the experiment on synthetic data, in real-world experiments, we see that the value of this loss is sufficiently low even when not directly minimizing it, possibly due to working in a high-dimensional embedding space where two random vectors are approximately orthogonal. A detailed discussion regarding $\mathcal{L}^{fso}$ is provided in Sections A.9 and A.10 with the conclusions presented in the main body of the article being mostly aligned with the conclusions that stem from detailed ablation studies.
After learning the directions, either with our method or with methods from the unsupervised state-of-the-art, we use the annotations available in the concept datasets to label the directions using Network Dissection (Bau et al., 2017). As described in Section 3.4, we make a basis change by projecting the network feature activations onto the learned decoding directions (for PCA and EDPP) or by directly utilizing the concept factors learned by NMF. While our approach is explicitly learning binary classifiers $\{{\bm{w}}_{i},b_{i}\}$ as concept detectors, PCA and NMF lack the estimation of an explicit classification threshold $b$. To form concept detectors in those cases, we follow Bau et al. (2017) and for each direction we choose $b$ to be equal to the top $k$-quantile among the projected activations sourced from the concept dataset. For NMF we use $k=0.005$, as suggested in Bau et al. (2017), while for PCA we use $k=0.2$ as suggested in Graziani et al. (2023a). In our experiments we also consider comparing with the clustering obtained using the natural latent space basis, i.e. the neuron directions, which in the tables we refer to as Natural. In that case, the classification threshold of the considered concept detectors is set to the top $0.005$ quantile of the feature activations in the direction of the neuron, as originally suggested in Network Dissection.
@@ 7.2 Faithfulness Assessment of the Encoding-Decoding Direction Pairs
In this section, we aim to evaluate the faithfulness of the encoding-decoding direction pairs, i.e. to assess whether the signal value extracted by the decoding direction is indeed encoded in the direction of the signal vector. Although we already validated that the direction pairs are faithful in the case of synthetic data (Section 6), in this paragraph, we aim to make the same evaluation in the case of real data, in the absence of ground truth. Based on activation maximization (Nguyen et al., 2019), we propose the following novel approach to address this evaluation.
@@
Figure 9: Our approach to assess the faithfulness of the direction pair: Starting from an image with the concept, we use deep dream to maximize the concept’s decoding direction and collect, during “dreaming”, the patch embeddings from the dream images at the regions with the concept. Subsequently, we fit a parametric line to those features in order to estimate the direction towards which they are moving when “dreaming”. We term this direction as the dreaming direction.
@@
Suppose that we evaluate the faithfulness of a particular direction pair $^{\bm{w}},\hat{{\bm{s}}}$ and let $b$ denote the respective concept detector’s bias. We use the concept detector (${\bm{w}},b$) in order to collect a set of images containing the concept. Subsequently, using these images as an initialization point, we use Deep Dream (Olah et al., 2017) to compute pre-images (Mahendran & Vedaldi, 2015; 2016), that is, images that maximize the decoding direction $\bm{w}$} at the spatial locations of where the concept was initially found. In this manner, we start from an image containing the concept and ask Deep Dream to amplify (in the input pixel space) whatever concept the detector can detect. We keep running the activation maximization loop for $K$ iterations and keep a record of how the features (at the locations with the concept) evolve while optimizing. Finally, we use all the recorded features collected during the “dreaming” optimization in order to fit a parametric line and estimate the direction that best describes feature evolvement while dreaming. We call the direction of this line the dreaming direction (Fig. 9). We conclude by comparing the dreaming direction of the concept for each image with the respective signal vector in terms of cosine similarity. In Figures 11 and 12 we provide histograms of these cosine similarities when applying our method to ResNet18 and EfficientNet. In all cases, even the ones without considering $\mathcal{L}^{fso}$, approximately $90\%$ of the signal vectors have cosine similarity with the dreaming directions above 0.7, indicating sufficient faithfulness. The effect of $\mathcal{L}^{fso}$ is to push the distribution to become more Gaussian. To validate the effectiveness of the sub-sampling strategy proposed in Section 4.3 we also provide histograms of cosine similarities between signal vectors and dreaming directions in the absence of sub-sampling. The results are depicted in Fig. 13 where, in that case, the deviation of the signal vectors from the dreaming direction is evident. Example dreaming pre-images for two cases are provided in Fig. 10 and further details and experiments may be found in Sections A.6 and A.9.
@@
Figure 14 depicts qualitative segmentation results obtained using our concept detectors. Visualizations are obtained using Network Dissection (Bau et al., 2017). In the figures, the Intersection over Union (IoU) scores refer to the whole validation split of the concept dataset and not individual image segmentations. We can verify that concept detectors appear to be mostly monosemantic. For each network architecture, we additionally report interpretability statistics provided by Network Dissection. In this report, a concept detector is considered to be interpretable whenever its IoU performance with the best possible concept in the concept dataset exceeds the threshold of $0.04$, which is equal to what was suggested in Network Dissection. Table 2 depicts this summary for ResNet18, while for the rest of the architectures, the results are summarized in Tables of the Appendix (Section A.12). More qualitative visualizations and visual comparisons with the unsupervised state-of-the-art are also included in the same Section of the Appendix.
Table 2: Network Dissection statistics for ResNet18 trained on Places365. For each concept category in Broden, we report two numbers: First, the number of concept detectors that were labeled with the name of a concept belonging to the category and second, the number of unique concept labels from the category that have been assigned to the set of the concept detectors. The table summarizes statistics for methods of the unsupervised state-of-the-art and for different values of the cluster count hyper-parameter II.
@@ II Method Color Object Part Material Scene Texture Total
As an exemplar interpretation of the report, Table 2 shows that when learning direction pairs for ResNet18 with $I=512$, Network Dissection reported that EDDP-C can identify 269 different concepts in the following categories: 66 objects, 165 scenes, 11 parts, 4 materials, 22 textures, and 1 color. The total number of interpretable concept detectors is 407. This preliminary statistical report together with the visualizations is given in order to provide an intuition regarding the interpretability of the concept detectors. A rigorous evaluation with comparisons is conducted in the Sections that follow.
@@ 7.4.1 Evaluation of Clustering Quality via Cluster Statistics
We report two statistical measures regarding the directional clustering. First, Coverage measures the percent of patches in the concept dataset that are classified positively by at least one of the concept detectors in the set, and second, Redundancy measures the average number of classifiers making a positive prediction for each patch in the dataset. Coverage may also be interpreted as the percent of pixels that have been assigned to at least one cluster, or the percent of pixels explained by the clustering, and redundancy may also be interpreted as the average number of clusters to which each patch belongs.
@@ 7.4.2 Evaluation of the Decoding Directions: Interpretability
Classification Let $\mathcal{T}_{i}$ denote the binary classification task of predicting whether a patch belongs to concept $i$. Given the concept label $l$ which was assigned to a concept detector from Network Dissection, we assess the detector’s performance on $\mathcal{T}_{l}$ using the standard binary classification metrics of Precision, Recall, F1, and Average Precision (AP). For each metric, we report average and standard deviation scores by aggregating across all detectors in the learned set.
Segmentation We additionally consider the performance of concept detectors in the binary semantic segmentation task and employ two metrics from Doumanoglou et al. (2023). Specifically, let $\phi_{i}(c,\mathcal{K})$ the Intersection Over Union for concept detector $i$ in identifying concept $c$ within the dataset $\mathcal{K}$. Define $c^{\star}_{i}=\text{argmax}_{c}\phi_{i}(c,\mathcal{K}_{t})$, indicating the concept label detected best by concept detector $i$ within the training subset of the dataset ($\mathcal{K}_{t}$). With $\mathcal{K}_{v}$ as the validation subset, we use the following interpretability scores $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$:
@@
The first metric $\mathcal{S}^{1}$ counts concept detectors with an IoU performance that exceeds a score threshold $\xi$. The second metric $\mathcal{S}^{2}$ uses the cardinality of the set $|.|$ to count the unique concept labels detected by the concept detectors with IoU above $\xi$. Both metrics become threshold-agnostic, by integrating on all $\xi\in[0,1]$. Viewing it from a different angle, $\mathcal{S}^{1}$ pertains to the segmentation efficiency of each concept detector individually, whereas $\mathcal{S}^{2}$ concerns the clustering diversity, meaning the capability of the detector ensemble to recognize a wide array of distinct concepts. Finally, we also report the standard mean Intersection over Union (mIoU) score, which is the average IoU performance score aggregated across all the detectors in the set.
Monosemanticity Even though Precision is a natural classification metric that can capture the monosemanticity of the clustering, it has the drawback of relying on the explicit labels that are available in the concept dataset. To overcome this labeling limitation, and inspired by Dreyer et al. (2024b), we additionally quantify the monosemanticity of the concepts identified by the concept detectors by utilizing the embedding space of CLIP ViT/B-16 (Radford et al., 2021). In particular, for each concept detector, we pick images from the validation split of the concept dataset for which the detector is most confident about the presence of the concept. We pick 100 unique images from the top confident predictions. Subsequently, we crop a rectangle around the area of the concept and obtain the CLIP embedding for this crop. Let $k^{i}_{1}$ and $k^{i}_{2}$ denote image indices in the set of selected images with $k^{i}_{1},k^{i}_{2}\in\{0,1,...99\}$ and $i$ denote the $i$-th concept detector. We use the following monosemanticity metric, which measures the average distance between CLIP embeddings (${\bm{e}}_{k^{i}_{1}},{\bm{e}}_{k^{i}_{2}}$) within a cluster, aggregated over all clusters:

### v2

@@ 6.2 Evaluation of the Encoding-Decoding Direction Pairs
In this experiment, signal vectors are jointly learned with the concept detectors, employing every proposed facet of the method presented in Section 4. Concept detectors are evaluated for their ability to detect each one of the ground-truth concepts, while signal vectors are evaluated for their alignment with the ground-truth concept encoding directions.
Decoding Directions (Concept Detectors): Using the synthetic pixel dataset $\{{\bm{x}}_{{\bm{p}}_{1}}\}\cup\{{\bm{x}}_{{\bm{p}}_{2}}\}$, we assess the ability of each learned concept detector to detect each of the ground-truth concepts. Since the method is unsupervised, the name of the concept that each detector identifies is unknown. For this reason, we need to implement direction labeling (Section 3.4). In Table 2, we present the Intersection over Union scores for each detector against actual concept classes. Zero values indicate complete purity and no mixing of the concepts, while scores of 1.0 indicate perfect concept detection. As an additional step, we also examine how well the learned filters extract signal values from representations. Since distractors and the latent space bias $\bm{c}$} are not directly estimated by our method, we can only quantify the extracted signal values as deviations from the dataset’s average value (See also Section 5.1, (29) and Section A.9). The Root Mean Squared Error (RMSE) between these extracted values and the ground truth, after subtracting the mean signal value, is noted as $0.06$. We also verified that in this experiment, where the distractor components are independent of the concept content contained in the features, the learned filters were orthogonal to all the vectors in the distractor basis $\bm{D}$}, without explicitly enforcing this to occur.
Encoding Directions: Fig. 8 - orange indicates perfect alignment between the learned signal vectors and the ground-truth concept encoding directions. We stress the fact that, in this experiment, signal vectors are driven by signal values extracted by the concept detectors, and this is different from Section 6.1, where signal vectors were tested under ideal input conditions.
@@ 6.3 Ablation Study
We also consider a case study where we learn the direction pairs by omitting $\mathcal{L}^{fso}$ from the objective function, i.e., learning the direction pairs without (/wo) it. In that case, the concept detectors still discovered the ground-truth directional clusters, as in Table 2. However, the decoding directions were less reliable in extracting the ground-truth signal values. The RMSE between the extracted values and the ground truth, after subtracting the mean signal value, was found to be $0.31$, which is substantially inferior to the standard case discussed above. The latter has an impact on the fidelity of the signal vectors. Fig. 8 purple visualizes the cosine similarity between the learned signal vectors and the ground-truth encoding directions for this case. The experiment shows that the signal direction estimation is less effective compared to when $\mathcal{L}^{fso}$ is taken into account.
6.4 Evaluation against other Baselines
The ground-truth encoding directions of this example (Table 15 in Section A.8) contain both positive and negative components, and the relationship among the encoding directions (and the distractors) is, in general, not orthogonal. In theory, and without the need for practical experiments, this example cannot be addressed by NMF, K-Means, or PCA. NMF would produce a signal basis of non-negative components. Similarly, the cluster centers of K-Means would point toward the positive quartile where the centroids reside. Finally, the non-orthogonal nature of the ground truth encoding directions implies that the PCA’s solution space is insufficient.
@@ 7 Experiments on Deep Image Classifiers
In this Section, we apply our method in real-world scenarios. In the following sections, we a) assess the faithfulness of the learned encoding-decoding direction pairs, b) provide qualitative segmentation results obtained by utilizing our learned concept detectors, c) compare with unsupervised baselines in interpretability and influence terms, d) conduct an ablation study to quantify the contribution of each proposed component, e) compare with supervised direction learning, aiming to provide more insights on the interpretability and influence of the directions, f) provide application use cases in which the learned direction pairs assist in providing global, local, and counterfactual model explanations, g) provide an example application in model correction.
@@ 7.1 Direction Learning and Evaluation Protocol
We apply our Encoding-Decoding Direction Pairs (EDDP) on the last convolution layer of five different networks: ResNet18 (resnet) trained on Places365 (places365), ResNet50 (resnet) trained on Moments in Time (MiT) (MomentsInTime), as well as EfficientNet (b0) (efficientnet), Inception-v3 (inception), and VGG16 (vgg) trained on ImageNet (imagenet). For all the networks that we study, when learning the direction pairs, we use the Broden (NetworkDissection) concept dataset, except for ResNet50, for which we use Broden-Action BrodenAction. The concept datasets feature dense pixel annotations; Broden includes $1197$ concepts across 63K~63K images in $5$ concept categories (object, part, material, texture, color), while Broden-Action incorporates an additional action category with $210$ labels and 23K~23K more images. We emphasize that when learning the direction pairs, we do not make use of the annotations that complement the concept datasets.
@@
For baselines, we consider directional clusterings based on either PCA, NMF, or the Natural latent space basis (referred to as Natural). For our EDDP, we examine two variants that are differentiated based on the choice of feature manipulation strategy used in Uncertainty Region Alignment. We use EDDP-U to refer to UFM and EDDP-C to CFM. Unless stated otherwise, our Encoding-Decoding Direction Pairs use the Augmented Lagrangian Loss from Section 4.5, all losses from Sections 4.2 and 3.3 (except $\mathcal{L}^{s}$), along with either $\mathcal{L}^{uur}$ (EDDP-U) or $\mathcal{L}^{cur}$ and $\mathcal{L}^{fso}$ (EDDP-C) from Sections 4.3 and 4.4.
After learning the directions with any unsupervised method, we use the annotations available in the concept datasets to label the directions using Network Dissection (NetworkDissection). As described in Section 3.4, we make a basis change by projecting the network feature activations onto the learned decoding directions (for PCA and EDPP) or by directly utilizing the concept factors learned by NMF. More details on the experimental setup for all methods, including baselines, are provided in Section A.12.
@@ 7.2 Faithfulness Assessment of the Encoding-Decoding Direction Pairs
In this Section, we aim to evaluate the faithfulness of the encoding-decoding direction pairs, i.e. to assess whether the signal value extracted by the decoding direction is indeed encoded in the direction of the signal vector. Although this was validated in synthetic data (Section 6), in this paragraph, we make the same evaluation in the case of real data, in the absence of ground truth. Based on activation maximization (feature-viz-survey), we propose the following novel approach to address this evaluation.
@@
Figure 9: Our approach to assess the faithfulness of the direction pair: Starting from an image with the concept, we use deep dream to maximize the concept’s decoding direction and collect the patch embeddings from the dream images at the regions with the concept. Subsequently, we fit a parametric line to those features in order to estimate the direction towards which they are moving when “dreaming”. We term this direction as the dreaming direction.
@@
Suppose that we evaluate the faithfulness of a particular direction pair $^{\bm{w}},\hat{{\bm{s}}}$ and let $b$ denote the respective concept detector’s bias. We use the concept detector (${\bm{w}},b$) to collect a set of images containing the concept. Subsequently, with these images as an initialization point, we use Deep Dream (feature-viz-olah) to compute pre-images (pre-image; pre-image2), that is, images that maximize the decoding direction $\bm{w}$} at the spatial locations in which the concept was initially found. In this manner, we start from an image containing the concept and ask Deep Dream to amplify (in the input pixel space) whatever concept the detector can detect. We keep running the activation maximization loop for $K$ iterations and keep a record of how the features (at the locations with the concept) evolve while optimizing. Finally, we use all the recorded features collected during the “dreaming” optimization to fit a parametric line and estimate the direction that best describes feature evolvement while dreaming. We call the direction of this line the dreaming direction (Fig. 9). We conclude by comparing the dreaming direction of the concept for each image with the respective signal vector in terms of cosine similarity.
In Figures 11 and 12, we provide histograms of these cosine similarities when applying our method to ResNet18 and EfficientNet. In all cases, even those without considering $\mathcal{L}^{fso}$, approximately $90\%$ of the signal vectors have cosine similarity with the dreaming directions above 0.7, indicating sufficient faithfulness. The effect of $\mathcal{L}^{fso}$ is to push the distribution to become more Gaussian. To validate the effectiveness of the sub-sampling strategy proposed in Section 4.3, we also provide histograms of cosine similarities between signal vectors and dreaming directions in the absence of sub-sampling. The results are depicted in Fig. 13 where, in that case, the deviation of the signal vectors from the dreaming direction is evident. Example dreaming pre-images for two cases are provided in Fig. 10 and further details and experiments may be found in Sections A.10 and A.11.
@@
Figure 14 presents qualitative segmentation results obtained using our concept detectors, confirming that they are largely monosemantic. These visualizations rely on Network Dissection to generate segmentation maps and calculate the Intersection over Union (IoU) scores, which represent the detector’s performance across the entire concept dataset’s validation split. A detector qualifies as interpretable if its IoU score surpasses the predefined threshold of $0.04$. We provide detailed quantitative interpretability statistics for each network architecture, exemplified by Table 3 for ResNet18 and other Tables in Section A.16. More extensive qualitative comparisons and statistics against unsupervised baselines are provided in the same Section of the Appendix.
Table 3: Network Dissection statistics for ResNet18 trained on Places365. For each concept category in Broden, we report two numbers: First, the number of concept detectors that were labeled with the name of a concept belonging to the category and second, the number of unique concept labels from the category that have been assigned to the set of the concept detectors. The table summarizes statistics for unsupervised methods and for different values of the cluster count hyper-parameter II.
@@ II Method Color Object Part Material Scene Texture Total
As an exemplary interpretation of the report, Table 3 shows that when learning direction pairs for ResNet18 with $I=512$, Network Dissection reported that EDDP-C can identify 269 different concepts in the following categories: 66 objects, 165 scenes, 11 parts, 4 materials, 22 textures, and 1 color. The total number of interpretable concept detectors is 407. This preliminary statistical report, together with the visualizations, is provided in order to offer intuition regarding the interpretability of the concept detectors. A rigorous evaluation with comparisons is conducted in the Sections that follow.
@@ 7.4.1 Evaluation of Clustering Quality via Cluster Statistics
We report two statistics for directional clustering. Coverage is the percentage of patches in the concept dataset that receive at least one positive prediction from the concept detectors. Equivalently, it is the percentage of pixels assigned to at least one cluster or explained by the clustering. Redundancy is the average number of detectors that classify each patch positively. Equivalently, it is the average number of clusters to which each patch belongs.
@@ 7.4.2 Evaluation of the Decoding Directions: Interpretability
Classification Let $\mathcal{T}_{i}$ denote the binary classification task of predicting whether a patch belongs to concept $i$. Given the concept label $l$, which was assigned to a concept detector with direction labeling, we assess the detector’s performance on $\mathcal{T}_{l}$ using the standard binary classification metrics of Precision, Recall, F1, and Average Precision (AP). For each metric, we report average and standard deviation scores by aggregating across all detectors in the learned set.
Segmentation We additionally consider the performance of concept detectors in the binary semantic segmentation task and employ two metrics from UIBE. Specifically, let $\phi_{i}(c,\mathcal{K})$ the Intersection Over Union for concept detector $i$ in identifying concept $c$ within the dataset $\mathcal{K}$. Define $c^{\star}_{i}=\text{argmax}_{c}\phi_{i}(c,\mathcal{K}_{t})$, indicating the concept label detected best by concept detector $i$ within the training subset of the dataset ($\mathcal{K}_{t}$). With $\mathcal{K}_{v}$ as the validation subset, we use the following interpretability scores $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$:
@@
The first metric $\mathcal{S}^{1}$ counts concept detectors with an IoU performance that exceeds a score threshold $\xi$. The second metric $\mathcal{S}^{2}$ uses the cardinality of the set $|.|$ to count the unique concept labels detected by the concept detectors with IoU above $\xi$. Both metrics become threshold-agnostic, by integrating over all $\xi\in[0,1]$. Viewing it from a different angle, $\mathcal{S}^{1}$ pertains to the segmentation efficiency of each concept detector individually, whereas $\mathcal{S}^{2}$ concerns the clustering diversity, meaning the capability of the detector ensemble to recognize a wide array of distinct concepts. Finally, we also report the standard mean Intersection over Union (mIoU) score, which is the average IoU performance score aggregated across all the detectors in the set.
Monosemanticity Even though Precision is a natural classification metric that can capture the monosemanticity of the clustering, it has the drawback of relying on the explicit labels that are available in the concept dataset. To overcome this labeling limitation, and inspired by pure-features, we additionally quantify the monosemanticity of the concepts identified by the concept detectors by utilizing the embedding space of CLIP ViT/B-16 (CLIP). In particular, for each concept detector, we pick images from the validation split of the concept dataset for which the detector is most confident about the presence of the concept. We select 100 unique images from the top confident predictions. Subsequently, we crop a rectangle around the area of the concept and obtain the CLIP embedding for this crop. Let $k^{i}_{1}$ and $k^{i}_{2}$ denote image indices in the set of selected images, with $k^{i}_{1},k^{i}_{2}\in\{0,1,...99\}$ and $i$ denoting the $i$-th concept detector. We use the following monosemanticity metric, which measures the average distance between CLIP embeddings (${\bm{e}}_{k^{i}_{1}},{\bm{e}}_{k^{i}_{2}}$) within a cluster, aggregated over all clusters: